Can you help me set up the algebra equation for this word problem?
A stamp collection consists of 3, 8, and 15 cent stamps. The number of 8 cent stamps is one less than triple the number of 3 cent stamps. The number of 15 cent stamps is six less than the number of 8 cent stamps. The total value of all the stamps is $2.47. Find the number of 8 cent stamps in the collection.

Answer by scott8148(6628)

let t="number of 3 cent stamps, e="number of 8 cent stamps" and f="number of 15 cent stamps"

"The number of 8 cent stamps is one less than triple the number of 3 cent stamps" ___ e = 3 t - 1
___ adding 1 and dividing by 3 ___ (e + 1) / 3 = t

"The number of 15 cent stamps is six less than the number of 8 cent stamps" ___ f = e - 6

"The total value of all the stamps is $2.47" ___ working in cents ___ 3 t + 8 e + 15 f = 247

substituting ___ 3 ((e+1) / 3) + 8 e + 15 (e - 6) = 247
distributing ___ e + 1 + 8 e + 15 e - 90 = 247
___ 24 e = 336
___ e = 14